# Dependability of UAV-Based Networks and Computing Systems: A Survey

### v1

@@ Acknowledgements.
- Haber et al. (2019a) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2019a. A reliability-aware computation offloading solution via UAV-mounted cloudlets. In 2019 IEEE 8th International Conference on Cloud Networking (CloudNet) (Coimbra, Portugal). IEEE.
- Haber et al. (2019b) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2019b. A reliability-aware computation offloading solution via UAV-mounted cloudlets. In 2019 IEEE 8th International Conference on Cloud Networking (CloudNet) (Coimbra, Portugal). IEEE.
- Haber et al. (2021a) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2021a. UAV-aided ultra-reliable low-latency computation offloading in future IoT networks. IEEE Trans. Commun. 69, 10 (Oct. 2021), 6838–6851.
- Haber et al. (2021b) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2021b. UAV-aided ultra-reliable low-latency computation offloading in future IoT networks. IEEE Trans. Commun. 69, 10 (Oct. 2021), 6838–6851.
@@
- Hang et al. (2020a) Bingqing Hang, Biling Zhang, Li Wang, Jingjing Wang, Yong Ren, and Zhu Han. 2020a. A user association policy for UAV-aided time-varying vehicular networks with MEC. In 2020 IEEE Wireless Communications and Networking Conference (WCNC) (Seoul, Korea (South)). IEEE.
- Hang et al. (2020b) Bingqing Hang, Biling Zhang, Li Wang, Jingjing Wang, Yong Ren, and Zhu Han. 2020b. A user association policy for UAV-aided time-varying vehicular networks with MEC. In 2020 IEEE Wireless Communications and Networking Conference (WCNC) (Seoul, Korea (South)). IEEE.
@@
- He et al. (2023a) Saifei He, Ming Cheng, Yijin Pan, Min Lin, and Wei-Ping Zhu. 2023a. Distributed access and offloading scheme for multiple UAVs assisted MEC networks. In 2023 IEEE 98th Vehicular Technology Conference (VTC2023-Fall) (Hong Kong, Hong Kong). IEEE, 1–5.
- He et al. (2023b) Saifei He, Ming Cheng, Yijin Pan, Min Lin, and Wei-Ping Zhu. 2023b. Distributed access and offloading scheme for multiple UAVs assisted MEC networks. In 2023 IEEE 98th Vehicular Technology Conference (VTC2023-Fall) (Hong Kong, Hong Kong). IEEE, 1–5.
@@
- Hirota et al. (2020a) Yusuke Hirota, Ittetsu Taniguchi, and Takao Onoye. 2020a. Parallelization of local path planning for high reliable autonomous drones. In 2020 International SoC Design Conference (ISOCC) (Yeosu, Korea (South)). IEEE.
- Hirota et al. (2020b) Yusuke Hirota, Ittetsu Taniguchi, and Takao Onoye. 2020b. Parallelization of local path planning for high reliable autonomous drones. In 2020 International SoC Design Conference (ISOCC) (Yeosu, Korea (South)). IEEE.
@@
- Hou et al. (2020c) Xiangwang Hou, Zhiyuan Ren, Jingjing Wang, Shuya Zheng, Wenchi Cheng, and Hailin Zhang. 2020c. Distributed fog computing for latency and reliability guaranteed swarm of drones. IEEE Access 8 (2020), 7117–7130.
- Hou et al. (2020d) Xiangwang Hou, Zhiyuan Ren, Jingjing Wang, Shuya Zheng, Wenchi Cheng, and Hailin Zhang. 2020d. Distributed fog computing for latency and reliability guaranteed swarm of drones. IEEE Access 8 (2020), 7117–7130.
@@
- Hou et al. (2020b) Xiangwang Hou, Zhiyuan Ren, Jingjing Wang, Shuya Zheng, and Hailin Zhang. 2020b. Latency and reliability oriented collaborative optimization for multi-UAV aided mobile edge computing system. In IEEE INFOCOM 2020 - IEEE Conference on Computer Communications Workshops (INFOCOM WKSHPS) (Toronto, ON, Canada). IEEE.
@@
- Jeong et al. (2020a) Hyeok-June Jeong, Suh-Yong Choi, Sung-Su Jang, and Young-Guk Ha. 2020a. Probability machine-learning-based communication and operation optimization for cloud-based UAVs. J. Supercomput. 76, 10 (Oct. 2020), 8101–8117.
- Jeong et al. (2020b) Hyeok-June Jeong, Suh-Yong Choi, Sung-Su Jang, and Young-Guk Ha. 2020b. Probability machine-learning-based communication and operation optimization for cloud-based UAVs. J. Supercomput. 76, 10 (Oct. 2020), 8101–8117.
@@
- Ju et al. (2024a) Tao Ju, Linjuan Li, Shuai Liu, and Yu Zhang. 2024a. A multi-UAV assisted task offloading and path optimization for mobile edge computing via multi-agent deep reinforcement learning. J. Netw. Comput. Appl. 229, 103919 (Sept. 2024), 103919.
- Ju et al. (2024b) Tao Ju, Linjuan Li, Shuai Liu, and Yu Zhang. 2024b. A multi-UAV assisted task offloading and path optimization for mobile edge computing via multi-agent deep reinforcement learning. J. Netw. Comput. Appl. 229, 103919 (Sept. 2024), 103919.
@@
- Jung et al. (2023b) Sooyeob Jung, Seongah Jeong, Jinkyu Kang, and Joonhyuk Kang. 2023b. Marine IoT systems with space–air–sea integrated networks: Hybrid LEO and UAV edge computing. IEEE Internet Things J. 10, 23 (Dec. 2023), 20498–20510.
- Jung et al. (2023c) Soyi Jung, Jae-Hyun Kim, David Mohaisen, and Joongheon Kim. 2023c. Truthful and performance-optimal computation outsourcing for aerial surveillance platforms via learning-based auction. Comput. Netw. 225, 109651 (April 2023), 109651.
@@
- Kaleem et al. (2019a) Zeeshan Kaleem, Muhammad Yousaf, Aamir Qamar, Ayaz Ahmad, Trung Q Duong, Wan Choi, and Abbas Jamalipour. 2019a. UAV-empowered disaster-resilient edge architecture for delay-sensitive communication. IEEE Netw. 33, 6 (Nov. 2019), 124–132.
- Kaleem et al. (2019b) Zeeshan Kaleem, Muhammad Yousaf, Aamir Qamar, Ayaz Ahmad, Trung Q Duong, Wan Choi, and Abbas Jamalipour. 2019b. UAV-empowered disaster-resilient edge architecture for delay-sensitive communication. IEEE Netw. 33, 6 (Nov. 2019), 124–132.
- Karam et al. (2024a) Sana Nasim Karam, Kashif Bilal, Junaid Shuja, Latif U Khan, Muhammad Bilal, and Muhammad Khurram Khan. 2024a. Intelligent IoT- and UAV-Assisted Architecture for Pipeline Monitoring in OGI. IT Prof. 26, 3 (May 2024), 46–54.
- Karam et al. (2024b) Sana Nasim Karam, Kashif Bilal, Junaid Shuja, Latif U Khan, Muhammad Bilal, and Muhammad Khurram Khan. 2024b. Intelligent IoT- and UAV-Assisted Architecture for Pipeline Monitoring in OGI. IT Prof. 26, 3 (May 2024), 46–54.
@@
- Li et al. (2024f) Bin Li, Rongrong Yang, Lei Liu, Junyi Wang, Ning Zhang, and Mianxiong Dong. 2024f. Robust Computation Offloading and Trajectory Optimization for Multi-UAV-Assisted MEC: A Multiagent DRL Approach. IEEE Internet Things J. 11, 3 (Feb. 2024), 4775–4786.
@@
- Li et al. (2020d) Meng Li, F Richard Yu, Pengbo Si, Ruizhe Yang, Zhuwei Wang, and Yanhua Zhang. 2020d. UAV-assisted data transmission in blockchain-enabled M2M communications with mobile edge computing. IEEE Netw. 34, 6 (Nov. 2020), 242–249.
@@
- Lin et al. (2021a) Xinhua Lin, Jing Zhang, Lin Xiang, and Xiaohu Ge. 2021a. Energy consumption optimization for UAV assisted private blockchain-based IIoT networks. In 2021 IEEE 94th Vehicular Technology Conference (VTC2021-Fall) (Norman, OK, USA). IEEE.
- Lin et al. (2021b) Xinhua Lin, Jing Zhang, Lin Xiang, and Xiaohu Ge. 2021b. Energy consumption optimization for UAV assisted private blockchain-based IIoT networks. In 2021 IEEE 94th Vehicular Technology Conference (VTC2021-Fall) (Norman, OK, USA). IEEE.
@@
- Liu et al. (2024d) Shuai Liu, Helin Yang, Liang Xiao, Mengting Zheng, Huabing Lu, and Zehui Xiong. 2024d. Learning-based resource management optimization for UAV-assisted MEC against jamming. IEEE Trans. Commun. 72, 8 (Aug. 2024), 4873–4886.
@@
- Liu et al. (2024c) Yinan Liu, Chao Yang, Xin Chen, and Fengyan Wu. 2024c. Joint hybrid caching and replacement scheme for UAV-assisted vehicular edge computing networks. IEEE Trans. Intell. Veh. 9, 1 (Jan. 2024), 866–878.
@@
- Lorincz et al. (2021a) Josip Lorincz, Adnan Tahirovic, and Biljana Risteska Stojkoska. 2021a. A novel real-time unmanned aerial vehicles-based disaster management framework. In 2021 29th Telecommunications Forum (TELFOR) (Belgrade, Serbia). IEEE.
- Lorincz et al. (2021b) Josip Lorincz, Adnan Tahirovic, and Biljana Risteska Stojkoska. 2021b. A novel real-time unmanned aerial vehicles-based disaster management framework. In 2021 29th Telecommunications Forum (TELFOR) (Belgrade, Serbia). IEEE.
@@
- Luo et al. (2021a) Yizhe Luo, Wenrui Ding, and Baochang Zhang. 2021a. Optimization of task scheduling and dynamic service strategy for multi-UAV-enabled mobile-edge computing system. IEEE Trans. Cogn. Commun. Netw. 7, 3 (Sept. 2021), 970–984.
- Luo et al. (2021b) Yizhe Luo, Wenrui Ding, and Baochang Zhang. 2021b. Optimization of task scheduling and dynamic service strategy for multi-UAV-enabled mobile-edge computing system. IEEE Trans. Cogn. Commun. Netw. 7, 3 (Sept. 2021), 970–984.
@@
- Ma and Sun (2022a) Zhi Ma and Songlin Sun. 2022a. Research on vehicle-to-road collaboration and end-to-end collaboration for multimedia services in the internet of vehicles. IEEE Access 10 (2022), 18146–18155.
- Ma and Sun (2022b) Zhi Ma and Songlin Sun. 2022b. Research on vehicle-to-road collaboration and end-to-end collaboration for multimedia services in the internet of vehicles. IEEE Access 10 (2022), 18146–18155.
@@
- Machida and Andrade (2021a) Fumio Machida and Ermeson Andrade. 2021a. PA-offload: Performability-aware adaptive fog offloading for drone image processing. In 2021 IEEE 5th International Conference on Fog and Edge Computing (ICFEC) (Melbourne, Australia). IEEE.
- Machida and Andrade (2021b) Fumio Machida and Ermeson Andrade. 2021b. PA-offload: Performability-aware adaptive fog offloading for drone image processing. In 2021 IEEE 5th International Conference on Fog and Edge Computing (ICFEC) (Melbourne, Australia). IEEE.
- Machida et al. (2023a) Fumio Machida, Qingyang Zhang, and Ermeson Andrade. 2023a. Performability analysis of adaptive drone computation offloading with fog computing. Future Gener. Comput. Syst. 145 (Aug. 2023), 121–135.
- Machida et al. (2023b) Fumio Machida, Qingyang Zhang, and Ermeson Andrade. 2023b. Performability analysis of adaptive drone computation offloading with fog computing. Future Gener. Comput. Syst. 145 (Aug. 2023), 121–135.
@@
- Manzini et al. (2023a) Thomas Manzini, Robin Murphy, David Merrick, and Justin Adams. 2023a. Wireless network demands of data products from small uncrewed aerial systems at hurricane Ian. In 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS) (Detroit, MI, USA). IEEE, 9941–9946.
- Manzini et al. (2023b) Thomas Manzini, Robin Murphy, David Merrick, and Justin Adams. 2023b. Wireless network demands of data products from small uncrewed aerial systems at hurricane Ian. In 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS) (Detroit, MI, USA). IEEE, 9941–9946.
@@
- Martínez-Sánchez et al. (2023) J Martínez-Sánchez, D Conde Morales, S Novoa, and P Arias. 2023. An Unmanned Aerial System for on-the-edge traffic monitoring in work-zone approach area. ISPRS - Int. Arch. Photogramm. Remote Sens. Spat. Inf. Sci. XLVIII-1/W2-2023 (Dec. 2023), 605–610.
@@
- Navardi and Mohsenin (2023a) Mozhgan Navardi and Tinoosh Mohsenin. 2023a. MLAE2: Metareasoning for latency-aware energy-efficient autonomous nano-drones. In 2023 IEEE International Symposium on Circuits and Systems (ISCAS) (Monterey, CA, USA). IEEE, 1–5.
- Navardi and Mohsenin (2023b) Mozhgan Navardi and Tinoosh Mohsenin. 2023b. MLAE2: Metareasoning for latency-aware energy-efficient autonomous nano-drones. In 2023 IEEE International Symposium on Circuits and Systems (ISCAS) (Monterey, CA, USA). IEEE, 1–5.
- Ngoenriang et al. (2022a) Napat Ngoenriang, Stephen John Turner, Dusit Niyato, and Sucha Supittayapornpong. 2022a. Joint UAV placement and data delivery in aerial inspection under uncertainties. IEEE Internet Things J. 9, 9 (May 2022), 6389–6403.
- Ngoenriang et al. (2022b) Napat Ngoenriang, Stephen John Turner, Dusit Niyato, and Sucha Supittayapornpong. 2022b. Joint UAV placement and data delivery in aerial inspection under uncertainties. IEEE Internet Things J. 9, 9 (May 2022), 6389–6403.
@@
- Ozger et al. (2018a) Mustafa Ozger, Michal Vondra, and Cicek Cavdar. 2018a. Towards beyond visual line of sight piloting of UAVs with ultra reliable low latency communication. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Ozger et al. (2018b) Mustafa Ozger, Michal Vondra, and Cicek Cavdar. 2018b. Towards beyond visual line of sight piloting of UAVs with ultra reliable low latency communication. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Pacheco et al. (2021a) Lucas Pacheco, Helder Oliveira, Denis Rosario, Zhongliang Zhao, Eduardo Cerqueira, Torsten Braun, and Paulo Mendes. 2021a. Towards the future of edge computing in the sky: Outlook and future directions. In 2021 17th International Conference on Distributed Computing in Sensor Systems (DCOSS) (Pafos, Cyprus). IEEE.
- Pacheco et al. (2021b) Lucas Pacheco, Helder Oliveira, Denis Rosario, Zhongliang Zhao, Eduardo Cerqueira, Torsten Braun, and Paulo Mendes. 2021b. Towards the future of edge computing in the sky: Outlook and future directions. In 2021 17th International Conference on Distributed Computing in Sensor Systems (DCOSS) (Pafos, Cyprus). IEEE.
@@
- Patnayak et al. (2021a) Chinmaya Patnayak, James E McClure, and Ryan K Williams. 2021a. WASP: A wearable super-computing platform for distributed intelligence in multi-agent systems. In 2021 IEEE High Performance Extreme Computing Conference (HPEC) (Waltham, MA, USA). IEEE.
- Patnayak et al. (2021b) Chinmaya Patnayak, James E McClure, and Ryan K Williams. 2021b. WASP: A wearable super-computing platform for distributed intelligence in multi-agent systems. In 2021 IEEE High Performance Extreme Computing Conference (HPEC) (Waltham, MA, USA). IEEE.
@@
- Picano and Fantacci (2024a) Benedetta Picano and Romano Fantacci. 2024a. A combined stochastic network calculus and matching theory approach for computational offloading in a heterogenous MEC environment. IEEE Trans. Netw. Serv. Manag. 21, 2 (April 2024), 1958–1968.
- Picano and Fantacci (2024b) Benedetta Picano and Romano Fantacci. 2024b. A combined stochastic network calculus and matching theory approach for computational offloading in a heterogenous MEC environment. IEEE Trans. Netw. Serv. Manag. 21, 2 (April 2024), 1958–1968.
@@
- Qureshi et al. (2024a) Khalid Ibrahim Qureshi, Lei Wang, Xuanrui Xiong, and Muhammad Ali Lodhi. 2024a. Asynchronous federated learning for resource allocation in software-defined internet of UAVs. IEEE Internet Things J. 11, 12 (June 2024), 20899–20911.
- Qureshi et al. (2024b) Khalid Ibrahim Qureshi, Lei Wang, Xuanrui Xiong, and Muhammad Ali Lodhi. 2024b. Asynchronous federated learning for resource allocation in software-defined internet of UAVs. IEEE Internet Things J. 11, 12 (June 2024), 20899–20911.
@@
- Seisa et al. (2022a) Achilleas Santi Seisa, Sumeet Gajanan Satpute, Bjorn Lindqvist, and George Nikolakopoulos. 2022a. An edge architecture oriented model predictive control scheme for an autonomous UAV mission. In 2022 IEEE 31st International Symposium on Industrial Electronics (ISIE) (Anchorage, AK, USA). IEEE.
- Seisa et al. (2022b) Achilleas Santi Seisa, Sumeet Gajanan Satpute, Bjorn Lindqvist, and George Nikolakopoulos. 2022b. An edge architecture oriented model predictive control scheme for an autonomous UAV mission. In 2022 IEEE 31st International Symposium on Industrial Electronics (ISIE) (Anchorage, AK, USA). IEEE.
@@
- Shao et al. (2023a) Ziling Shao, Helin Yang, Liang Xiao, Wei Su, and Zehui Xiong. 2023a. Energy and latency-aware resource management for UAV-assisted mobile edge computing against jamming. In GLOBECOM 2023 - 2023 IEEE Global Communications Conference (Kuala Lumpur, Malaysia). IEEE.
- Shao et al. (2023b) Ziling Shao, Helin Yang, Liang Xiao, Wei Su, and Zehui Xiong. 2023b. Energy and latency-aware resource management for UAV-assisted mobile edge computing against jamming. In GLOBECOM 2023 - 2023 IEEE Global Communications Conference (Kuala Lumpur, Malaysia). IEEE.
@@
- Sharma et al. (2019a) Vishal Sharma, Ilsun You, Dushantha Nalin K Jayakody, Daniel Gutierrez Reina, and Kim-Kwang Raymond Choo. 2019a. Neural-blockchain-based ultrareliable caching for edge-enabled UAV networks. IEEE Trans. Industr. Inform. 15, 10 (Oct. 2019), 5723–5736.
- Sharma et al. (2019b) Vishal Sharma, Ilsun You, Dushantha Nalin K Jayakody, Daniel Gutierrez Reina, and Kim-Kwang Raymond Choo. 2019b. Neural-blockchain-based ultrareliable caching for edge-enabled UAV networks. IEEE Trans. Industr. Inform. 15, 10 (Oct. 2019), 5723–5736.
@@
- Taleb et al. (2023a) Tarik Taleb, Nassim Sehad, Zinelaabidine Nadir, and Jaeseung Song. 2023a. VR-Based Immersive Service Management in B5G Mobile Systems: A UAV Command and Control Use Case. IEEE Internet Things J. 10, 6 (March 2023), 5349–5363.
- Taleb et al. (2023b) Tarik Taleb, Nassim Sehad, Zinelaabidine Nadir, and Jaeseung Song. 2023b. VR-Based Immersive Service Management in B5G Mobile Systems: A UAV Command and Control Use Case. IEEE Internet Things J. 10, 6 (March 2023), 5349–5363.
@@
- Thapliyal et al. (2024a) Shardul Thapliyal, Rajoo Pandey, and Chhagan Charan. 2024a. Energy-harvesting-enabled UAV relay-assisted NOMA network under finite blocklength. IEEE Internet Things J. 11, 13 (July 2024), 23255–23267.
- Thapliyal et al. (2024b) Shardul Thapliyal, Rajoo Pandey, and Chhagan Charan. 2024b. Energy-harvesting-enabled UAV relay-assisted NOMA network under finite blocklength. IEEE Internet Things J. 11, 13 (July 2024), 23255–23267.
@@
- Tripathi et al. (2023a) Shraddha Tripathi, Om Jee Pandey, Linga Reddy Cenkeramaddi, and Rajesh M Hegde. 2023a. A socially-aware radio map framework for improving QoS of UAV-assisted MEC networks. IEEE Trans. Netw. Serv. Manag. 20, 1 (March 2023), 342–356.
- Tripathi et al. (2023b) Shraddha Tripathi, Om Jee Pandey, Linga Reddy Cenkeramaddi, and Rajesh M Hegde. 2023b. A socially-aware radio map framework for improving QoS of UAV-assisted MEC networks. IEEE Trans. Netw. Serv. Manag. 20, 1 (March 2023), 342–356.
@@
- Venkatagiri et al. (2018a) Radha Venkatagiri, Karthik Swaminathan, Chung-Ching Lin, Liang Wang, Alper Buyuktosunoglu, Pradip Bose, and Sarita Adve. 2018a. Impact of software approximations on the resiliency of a video summarization system. In 2018 48th Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN) (Luxembourg City). IEEE.
- Venkatagiri et al. (2018b) Radha Venkatagiri, Karthik Swaminathan, Chung-Ching Lin, Liang Wang, Alper Buyuktosunoglu, Pradip Bose, and Sarita Adve. 2018b. Impact of software approximations on the resiliency of a video summarization system. In 2018 48th Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN) (Luxembourg City). IEEE.
@@
- Wang et al. (2020e) Guangying Wang, Xiangbin Yu, Fangcheng Xu, and Jiali Cai. 2020e. Task offloading and resource allocation for UAV-assisted mobile edge computing with imperfect channel estimation over Rician fading channels. EURASIP J. Wirel. Commun. Netw. 2020, 1 (Dec. 2020).

### v2

@@ Acknowledgements.
- Haber et al. (2019) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2019. A reliability-aware computation offloading solution via UAV-mounted cloudlets. In 2019 IEEE 8th International Conference on Cloud Networking (CloudNet) (Coimbra, Portugal). IEEE.
- Haber et al. (2021) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2021. UAV-aided ultra-reliable low-latency computation offloading in future IoT networks. IEEE Trans. Commun. 69, 10 (Oct. 2021), 6838–6851.
@@
- Hang et al. (2020) Bingqing Hang, Biling Zhang, Li Wang, Jingjing Wang, Yong Ren, and Zhu Han. 2020. A user association policy for UAV-aided time-varying vehicular networks with MEC. In 2020 IEEE Wireless Communications and Networking Conference (WCNC) (Seoul, Korea (South)). IEEE.
@@
- He et al. (2023) Saifei He, Ming Cheng, Yijin Pan, Min Lin, and Wei-Ping Zhu. 2023. Distributed access and offloading scheme for multiple UAVs assisted MEC networks. In 2023 IEEE 98th Vehicular Technology Conference (VTC2023-Fall) (Hong Kong, Hong Kong). IEEE, 1–5.
@@
- Hirota et al. (2020) Yusuke Hirota, Ittetsu Taniguchi, and Takao Onoye. 2020. Parallelization of local path planning for high reliable autonomous drones. In 2020 International SoC Design Conference (ISOCC) (Yeosu, Korea (South)). IEEE.
@@
- Hou et al. (2020b) Xiangwang Hou, Zhiyuan Ren, Jingjing Wang, Shuya Zheng, Wenchi Cheng, and Hailin Zhang. 2020b. Distributed fog computing for latency and reliability guaranteed swarm of drones. IEEE Access 8 (2020), 7117–7130.
@@
- Jeong et al. (2020) Hyeok-June Jeong, Suh-Yong Choi, Sung-Su Jang, and Young-Guk Ha. 2020. Probability machine-learning-based communication and operation optimization for cloud-based UAVs. J. Supercomput. 76, 10 (Oct. 2020), 8101–8117.
@@
- Ju et al. (2024) Tao Ju, Linjuan Li, Shuai Liu, and Yu Zhang. 2024. A multi-UAV assisted task offloading and path optimization for mobile edge computing via multi-agent deep reinforcement learning. J. Netw. Comput. Appl. 229, 103919 (Sept. 2024), 103919.
@@
- Jung et al. (2023b) Soyi Jung, Jae-Hyun Kim, David Mohaisen, and Joongheon Kim. 2023b. Truthful and performance-optimal computation outsourcing for aerial surveillance platforms via learning-based auction. Comput. Netw. 225, 109651 (April 2023), 109651.
@@
- Kaleem et al. (2019) Zeeshan Kaleem, Muhammad Yousaf, Aamir Qamar, Ayaz Ahmad, Trung Q Duong, Wan Choi, and Abbas Jamalipour. 2019. UAV-empowered disaster-resilient edge architecture for delay-sensitive communication. IEEE Netw. 33, 6 (Nov. 2019), 124–132.
- Karam et al. (2024) Sana Nasim Karam, Kashif Bilal, Junaid Shuja, Latif U Khan, Muhammad Bilal, and Muhammad Khurram Khan. 2024. Intelligent IoT- and UAV-Assisted Architecture for Pipeline Monitoring in OGI. IT Prof. 26, 3 (May 2024), 46–54.
@@
- Lin et al. (2021) Xinhua Lin, Jing Zhang, Lin Xiang, and Xiaohu Ge. 2021. Energy consumption optimization for UAV assisted private blockchain-based IIoT networks. In 2021 IEEE 94th Vehicular Technology Conference (VTC2021-Fall) (Norman, OK, USA). IEEE.
@@
- Liu et al. (2024c) Shuai Liu, Helin Yang, Liang Xiao, Mengting Zheng, Huabing Lu, and Zehui Xiong. 2024c. Learning-based resource management optimization for UAV-assisted MEC against jamming. IEEE Trans. Commun. 72, 8 (Aug. 2024), 4873–4886.
@@
- Lorincz et al. (2021) Josip Lorincz, Adnan Tahirovic, and Biljana Risteska Stojkoska. 2021. A novel real-time unmanned aerial vehicles-based disaster management framework. In 2021 29th Telecommunications Forum (TELFOR) (Belgrade, Serbia). IEEE.
@@
- Luo et al. (2021) Yizhe Luo, Wenrui Ding, and Baochang Zhang. 2021. Optimization of task scheduling and dynamic service strategy for multi-UAV-enabled mobile-edge computing system. IEEE Trans. Cogn. Commun. Netw. 7, 3 (Sept. 2021), 970–984.
@@
- Ma and Sun (2022) Zhi Ma and Songlin Sun. 2022. Research on vehicle-to-road collaboration and end-to-end collaboration for multimedia services in the internet of vehicles. IEEE Access 10 (2022), 18146–18155.
@@
- Machida and Andrade (2021) Fumio Machida and Ermeson Andrade. 2021. PA-offload: Performability-aware adaptive fog offloading for drone image processing. In 2021 IEEE 5th International Conference on Fog and Edge Computing (ICFEC) (Melbourne, Australia). IEEE.
- Machida et al. (2023) Fumio Machida, Qingyang Zhang, and Ermeson Andrade. 2023. Performability analysis of adaptive drone computation offloading with fog computing. Future Gener. Comput. Syst. 145 (Aug. 2023), 121–135.
@@
- Manzini et al. (2023) Thomas Manzini, Robin Murphy, David Merrick, and Justin Adams. 2023. Wireless network demands of data products from small uncrewed aerial systems at hurricane Ian. In 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS) (Detroit, MI, USA). IEEE, 9941–9946.
@@
- Mart´ınez-Sánchez et al. (2023) J Martínez-Sánchez, D Conde Morales, S Novoa, and P Arias. 2023. An Unmanned Aerial System for on-the-edge traffic monitoring in work-zone approach area. ISPRS - Int. Arch. Photogramm. Remote Sens. Spat. Inf. Sci. XLVIII-1/W2-2023 (Dec. 2023), 605–610.
@@
- Navardi and Mohsenin (2023) Mozhgan Navardi and Tinoosh Mohsenin. 2023. MLAE2: Metareasoning for latency-aware energy-efficient autonomous nano-drones. In 2023 IEEE International Symposium on Circuits and Systems (ISCAS) (Monterey, CA, USA). IEEE, 1–5.
- Ngoenriang et al. (2022) Napat Ngoenriang, Stephen John Turner, Dusit Niyato, and Sucha Supittayapornpong. 2022. Joint UAV placement and data delivery in aerial inspection under uncertainties. IEEE Internet Things J. 9, 9 (May 2022), 6389–6403.
@@
- Ozger et al. (2018) Mustafa Ozger, Michal Vondra, and Cicek Cavdar. 2018. Towards beyond visual line of sight piloting of UAVs with ultra reliable low latency communication. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Pacheco et al. (2021) Lucas Pacheco, Helder Oliveira, Denis Rosario, Zhongliang Zhao, Eduardo Cerqueira, Torsten Braun, and Paulo Mendes. 2021. Towards the future of edge computing in the sky: Outlook and future directions. In 2021 17th International Conference on Distributed Computing in Sensor Systems (DCOSS) (Pafos, Cyprus). IEEE.
@@
- Patnayak et al. (2021) Chinmaya Patnayak, James E McClure, and Ryan K Williams. 2021. WASP: A wearable super-computing platform for distributed intelligence in multi-agent systems. In 2021 IEEE High Performance Extreme Computing Conference (HPEC) (Waltham, MA, USA). IEEE.
@@
- Picano and Fantacci (2024) Benedetta Picano and Romano Fantacci. 2024. A combined stochastic network calculus and matching theory approach for computational offloading in a heterogenous MEC environment. IEEE Trans. Netw. Serv. Manag. 21, 2 (April 2024), 1958–1968.
@@
- Qureshi et al. (2024) Khalid Ibrahim Qureshi, Lei Wang, Xuanrui Xiong, and Muhammad Ali Lodhi. 2024. Asynchronous federated learning for resource allocation in software-defined internet of UAVs. IEEE Internet Things J. 11, 12 (June 2024), 20899–20911.
@@
- Seisa et al. (2022) Achilleas Santi Seisa, Sumeet Gajanan Satpute, Bjorn Lindqvist, and George Nikolakopoulos. 2022. An edge architecture oriented model predictive control scheme for an autonomous UAV mission. In 2022 IEEE 31st International Symposium on Industrial Electronics (ISIE) (Anchorage, AK, USA). IEEE.
@@
- Shao et al. (2023) Ziling Shao, Helin Yang, Liang Xiao, Wei Su, and Zehui Xiong. 2023. Energy and latency-aware resource management for UAV-assisted mobile edge computing against jamming. In GLOBECOM 2023 - 2023 IEEE Global Communications Conference (Kuala Lumpur, Malaysia). IEEE.
@@
- Sharma et al. (2019) Vishal Sharma, Ilsun You, Dushantha Nalin K Jayakody, Daniel Gutierrez Reina, and Kim-Kwang Raymond Choo. 2019. Neural-blockchain-based ultrareliable caching for edge-enabled UAV networks. IEEE Trans. Industr. Inform. 15, 10 (Oct. 2019), 5723–5736.
@@
- Taleb et al. (2023) Tarik Taleb, Nassim Sehad, Zinelaabidine Nadir, and Jaeseung Song. 2023. VR-Based Immersive Service Management in B5G Mobile Systems: A UAV Command and Control Use Case. IEEE Internet Things J. 10, 6 (March 2023), 5349–5363.
@@
- Thapliyal et al. (2024) Shardul Thapliyal, Rajoo Pandey, and Chhagan Charan. 2024. Energy-harvesting-enabled UAV relay-assisted NOMA network under finite blocklength. IEEE Internet Things J. 11, 13 (July 2024), 23255–23267.
@@
- Tripathi et al. (2023) Shraddha Tripathi, Om Jee Pandey, Linga Reddy Cenkeramaddi, and Rajesh M Hegde. 2023. A socially-aware radio map framework for improving QoS of UAV-assisted MEC networks. IEEE Trans. Netw. Serv. Manag. 20, 1 (March 2023), 342–356.
@@
- Venkatagiri et al. (2018) Radha Venkatagiri, Karthik Swaminathan, Chung-Ching Lin, Liang Wang, Alper Buyuktosunoglu, Pradip Bose, and Sarita Adve. 2018. Impact of software approximations on the resiliency of a video summarization system. In 2018 48th Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN) (Luxembourg City). IEEE.
@@
- Wang et al. (2020d) Guangying Wang, Xiangbin Yu, Fangcheng Xu, and Jiali Cai. 2020d. Task offloading and resource allocation for UAV-assisted mobile edge computing with imperfect channel estimation over Rician fading channels. EURASIP J. Wirel. Commun. Netw. 2020, 1 (Dec. 2020).